# Online Conformal Probabilistic Numerics via Adaptive Edge-Cloud Offloading

### v1

@@ 5.1 Setup
At each round $t$, we randomly and independently generate a Hermitian matrix $A_{t}$ as $A_{t}=Q_{t}\Lambda_{t}Q_{t}$, where $n_{t}\times n_{t}$ orthogonal matrix $Q_{t}$ follows the Haar distribution and the diagonal matrix $\Lambda_{t}$ has diagonal elements drawn from a Gamma distribution with shape parameter $10$, scale parameter $1$, and location parameter 00. The dimension $n_{t}$ is drawn i.i.d. and uniformly from the interval $[500,1000]$. Furthermore, each element of vector $b_{t}$ is sampled i.i.d. from the standard multivariate Gaussian distribution. We set the target miscoverage rate $\alpha$ as $0.1$. The step size $\gamma$ and initial threshold $\lambda_{1}$ are set as $0.05$ and $0.95$, respectively. For the hyperparameter $\theta$, we use grid search to find a suitable one as $-5.5$.
@@
  • Time-varying edge computing budget: In this more challenging scenario, the number of allowed iterations starts at $I_{t}=0.1n_{t}$ for $t=1$ to $t=1500$, then it decreases to $I_{t}=0.05n_{t}$ for $t=1500$ to $t=3500$, and finally it increases to $I_{t}=0.15n_{t}$ for $t=3500$ to $t=5000$.
@@ 5.2 Baselines
  • PLS: Conventional PLS directly adopts the $(1-\alpha)$ uncertainty set as in (2.2).
@@ 5.3 Performance Metrics
As performance measures, we report empirical coverage, empirical uncertainty set size, and empirical cloud-to-edge feedback rate evaluated on a sequence $\{A_{t},b_{t}\}_{t=1}^{T}$ of duration $T=5000$. The empirical coverage at round $t$ measures the fraction of times in which the exact solution $x^{*}_{\tau}$ was included in the corresponding uncertainty set $\mathcal{C}_{t}$ from time $\tau=1$ until time $\tau=t$, i.e.,
@@
The empirical uncertainty set size at time step $t$ is defined as
empirical set $\text{empirical set size}=\frac{1}{t}\sum_{\tau=1}^{t}|\mathcal{C}_{\tau}|.$ (28)
@@ 5.4 Results
Fig. 3(a) demonstrates that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback guarantee the long-term coverage condition (2), thereby validating Proposition 1. In contrast, confirming prior art (Hegde et al., 2024; Wenger and Hennig, 2020), PLS produces an overly conservative uncertainty set that always covers the true solution. PLS achieves coverage of $1$ by yielding much larger uncertainty sets than both OCP-PLS and OCP-PLS with full cloud-to-edge feedback, as shown in Fig. 3(b). Furthermore, as illustrated in (2), OCP-PLS can reduce the cloud-to-edge feedback rate by approximately $40\%$ as compared to OCP-PLS with full cloud-to-edge feedback, while producing sets with similar sizes.
Consider now the setting with a time-varying computing budget at the edge. Fig. 4 marks the times at which the computing budget changes with vertical dashed lines. Fig. 4(a) verifies that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback maintain the guaranteed long-term coverage condition proved in Proposition 1 even under time-varying computing budgets. As in the previous example, PLS provides perfect coverage at the cost of excessively large uncertainty sets. In contrast, as shown in Fig. 4(b), OCP-PLS responds naturally to changes in edge computing budget, as the set sizes increase during periods of reduced computing budget ($t=1500$ to $t=3500$), while decreasing when the computing budget is higher ($t=3500$ to $t=5000$). Throughout these variations, OCP-PLS consistently maintains a significantly lower cloud-to-edge feedback rate than OCP-PLS with full cloud-to-edge feedback, as shown in Fig. 4(c), adapting the feedback rate to the varying computing budgets. This demonstrates OCP-PLS’s ability to balance reliability and communication efficiency under dynamic edge computing budgets.
@@ 6 Conclusion
An important challenge in edge-cloud computing architectures is to provide users with timely and reliable responses under limited computational budgets. This work has made steps towards addressing this challenge by leveraging probabilistic numerics (PN) and online conformal prediction (OCP), an online calibration method previously studied in the context of predictive models. The proposed methodology, referred to as OCP-PLS calibrates the uncertainty sets produced by probabilistic linear solvers (PLS), ensuring provable coverage guarantees for the true solution. OCP-PLS supports an adaptive mechanism for the use of cloud resources that reduces the communication overhead between edge and cloud, as well as the computational load of the cloud processors. Experiments for tasks requiring the sequential solution of linear systems validate the theoretical reliability, as well as the efficiency, of OCP-PLS. Future work may investigate the generalization of the proposed approach to other numerical problems beyond linear systems (Wenger et al., 2022; Pförtner et al., 2022; Hegde et al., 2024).
Appendix: Extension to Delayed Cloud-to-Edge Feedback
@@
Note that Proposition 2 reduces to Proposition 1 in Section 4.3 by setting $Q=1$ and $d_{t}=0$, i.e., when there is no delay in the cloud-edge feedback.

### v2

@@ 5.1 Setup
At each round $t$, we randomly and independently generate a Hermitian matrix $A_{t}$ as $A_{t}=Q_{t}\Lambda_{t}Q_{t}$, where $n_{t}\times n_{t}$ orthogonal matrix $Q_{t}$ follows the Haar distribution and the diagonal matrix $\Lambda_{t}$ has diagonal elements drawn from a Gamma distribution with shape parameter $10$, scale parameter $1$, and location parameter 00. The dimension $n_{t}$ is drawn i.i.d. and uniformly from the interval $[500,1000]$. Furthermore, each element of vector $b_{t}$ is sampled i.i.d. from the standard multivariate Gaussian distribution. We use the prior $x\sim\mathcal{N}(0,I)$, and BayesCG for the search directions (Cockayne et al., 2019b).
We set the target miscoverage rate $\alpha$ as $0.1$. The step size $\gamma$ and initial threshold $\lambda_{1}$ are set as $0.05$ and $0.99$, respectively. For the hyperparameter $\theta$, we use grid search to find a suitable one as $-3.5$.
@@
  • Time-varying edge computing budget: In this more challenging scenario, the number of allowed iterations starts at $I_{t}=0.1n_{t}$ for $t=1$ to $t=1500$, then it decreases to $I_{t}=0.005n_{t}$ for $t=1500$ to $t=3500$, and finally it increases to $I_{t}=0.15n_{t}$ for $t=3500$ to $t=5000$.
@@ 5.2 Baselines
  • PLS: Conventional PLS directly adopts the $(1-\alpha)$-HPD set as in (2.2).
@@ 5.3 Performance Metrics
As performance measures, we report empirical coverage, empirical HPD set size, and empirical cloud-to-edge feedback rate evaluated on a sequence $\{A_{t},b_{t}\}_{t=1}^{T}$ of duration $T=5000$. The empirical coverage at round $t$ measures the fraction of times in which the exact solution $x^{*}_{\tau}$ was included in the corresponding HPD set $\mathcal{C}_{t}$ from time $\tau=1$ until time $\tau=t$, i.e.,
@@
The size of the HDP set is measured by the empirical volume radius as
empirical volume $\text{empirical volume radius}=\frac{1}{t}\sum_{\tau=1}^{t}|\mathcal{C}_{\tau}|^{1/r_{\tau}}.$ (29)
@@ 5.4 Results
Fig. 3(a) demonstrates that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback guarantee the long-term coverage condition (2), thereby validating Proposition 1. In contrast, confirming prior art (Hegde et al., 2024; Wenger and Hennig, 2020), PLS produces an overly conservative HPD set that always covers the true solution. PLS achieves coverage of $1$ by yielding much larger HPD sets than both OCP-PLS and OCP-PLS with full cloud-to-edge feedback, as shown in
Fig. 3(b). Furthermore, as illustrated in (2), OCP-PLS can reduce the cloud-to-edge feedback rate by approximately $40\%$ as compared to OCP-PLS with full cloud-to-edge feedback, while producing sets with similar sizes.
Consider now the setting with a time-varying computing budget at the edge. Fig. 4 marks the times at which the computing budget changes with vertical dashed lines. Fig. 4(a) verifies that both OCP-PLS and OCP-PLS with full cloud-to-edge feedback maintain the guaranteed long-term coverage condition proved in Proposition 1 even under time-varying computing budgets. As in the previous example, PLS provides perfect coverage at the cost of excessively large HPD sets. In contrast, as shown in Fig. 4(b), OCP-PLS responds naturally to changes in edge computing budget, as the set sizes increase during periods of reduced computing budget ($t=1500$ to $t=3500$), while decreasing when the computing budget is higher ($t=3500$ to $t=5000$). Throughout these variations, OCP-PLS consistently
maintains a significantly lower cloud-to-edge feedback rate than OCP-PLS with full cloud-to-edge feedback, as shown in Fig. 4(c), adapting the feedback rate to the varying computing budgets. This demonstrates OCP-PLS’s ability to balance reliability and communication efficiency under dynamic edge computing budgets.
@@ 6 Conclusion
An important challenge in edge-cloud computing architectures is to provide users with timely and reliable responses under limited computational budgets. This work has made steps towards addressing this challenge by leveraging probabilistic numerics (PN) and online conformal prediction (OCP), an online calibration method previously studied in the context of predictive models. The proposed methodology, referred to as OCP-PLS calibrates the HPD sets produced by probabilistic linear solvers (PLS), ensuring provable coverage guarantees for the true solution. OCP-PLS supports an adaptive mechanism for the use of cloud resources that reduces the communication overhead between edge and cloud, as well as the computational load of the cloud processors. Experiments for tasks requiring the sequential solution of linear systems validate the theoretical reliability, as well as the efficiency, of OCP-PLS. Future work may investigate the generalization of the proposed approach to other numerical problems beyond linear systems (Wenger et al., 2022; Pförtner et al., 2022; Hegde et al., 2024).
Appendix A: Extension to Delayed
Cloud-to-Edge Feedback
@@
Note that Proposition 2 reduces to Proposition 1 in Section 4.3 by setting $T^{\max}=1$ and $d_{t}=0$, i.e., when there is no delay in the cloud-edge feedback.